# Provable Benefit of Annealed Langevin Monte Carlo for Non-log-concave Sampling

### v1

@@ 6 Conclusions and Future Work
In this paper, we have explored the complexity of annealed LMC for sampling from a non-log-concave probability measure without relying on log-concavity or isoperimetry. Central to our analysis are the Girsanov theorem and optimal transport techniques, thus circumventing the need for isoperimetric inequalities. While our proof primarily focuses on the annealing scheme as described in Equation 3, there is potential for adaptation to more general interpolations. Further exploration of these applications will be a focus of our future research. Technically, our proof methodology could be expanded to include a broader range of target distributions beyond those with smooth potentials, such as those with Hölder-continuous gradients [chatterji2020langevin, erdogdu2021on, liang2022a, nguyen2023unadjusted, fan2023nestrov, fan2023improved], or even heavy-tailed target distributions [mousavi-hosseini2023towards, ye2024an]. Furthermore, eliminating the necessity of assuming global minimizers of the potential function would enhance the practical utility of our algorithm. Theoretically, proving Conjecture 1 or obtaining a polynomial bound of ${\cal A}_{\eta,\lambda}(\pi)$ under less restrictive conditions would further validate that our complexity bounds are genuinely polynomial. Finally, while this work emphasizes the upper bounds for non-log-concave sampling, an intriguing future avenue would be to verify the tightness of these bounds and explore potential lower bounds for this problem [chewi2022the, chewi2023query, chewi2023fisher].
[…]
Appendix A Proof of Lemma 3
Our proof is inspired by [liang2023a].
@@ Appendix A Proof of Lemma 3
We only consider the nontrivial case $\eta_{0}\in(0,1]$. The potential of $\pi_{0}$ is $V_{0}:=\eta_{0}V+\frac{\lambda_{0}}{2}\left\|\cdot\right\|^{2}$, which is $(\lambda_{0}-\eta_{0}\beta)$-strongly-convex and $(\lambda_{0}+\eta_{0}\beta)$-smooth. Note that for any fixed point $x^{\prime}\in\mathbb{R}^{d}$,
@@
$\pi_{0}(x)\propto\exp\left(-V_{0}(x)\right)\leq\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x-x^{\prime}\right\|^{2}\right).$
@@
The right-hand side is the unnormalized density of $\pi_{0}^{\prime}:=\operatorname{{\cal N}}\left(x^{\prime}-\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}-\eta_{0}\beta},\frac{1}{\lambda_{0}-\eta_{0}\beta}I\right)$. The rejection sampling algorithm is as follows: sample $X\sim\pi_{0}^{\prime}$, and accept $X$ as a sample from $\pi_{0}$ with probability
@@
$\exp\left(-V_{0}(X)+V_{0}(x^{\prime})+\left\langle\nabla V_{0}(x^{\prime}),X-x^{\prime}\right\rangle+\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|X-x^{\prime}\right\|^{2}\right)\in(0,1].$
@@
By [chewi2022log, Theorem 7.1.1], the number of queries to the oracle of $V$ until acceptance follows a geometric distribution with mean
@@
$\frac{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x-x^{\prime}\right\|^{2}\right){\rm d}x}{\int\exp\left(-V_{0}(x)\right){\rm d}x}$
@@
$\leq\frac{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x-x^{\prime}\right\|^{2}\right){\rm d}x}{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}+\eta_{0}\beta}{2}\left\|x-x^{\prime}\right\|^{2}\right){\rm d}x}$
@@
$=\frac{\int\exp\left(-\left\langle\nabla V_{0}(x^{\prime}),x\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x\right\|^{2}\right){\rm d}x}{\int\exp\left(-\left\langle\nabla V_{0}(x^{\prime}),x\right\rangle-\frac{\lambda_{0}+\eta_{0}\beta}{2}\left\|x\right\|^{2}\right){\rm d}x}$
@@
$=\frac{\int\exp\left(-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x+\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}-\eta_{0}\beta}\right\|^{2}+\frac{\left\|\nabla V_{0}(x^{\prime})\right\|^{2}}{2(\lambda_{0}-\eta_{0}\beta)}\right){\rm d}x}{\int\exp\left(-\frac{\lambda_{0}+\eta_{0}\beta}{2}\left\|x+\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}+\eta_{0}\beta}\right\|^{2}+\frac{\left\|\nabla V_{0}(x^{\prime})\right\|^{2}}{2(\lambda_{0}+\eta_{0}\beta)}\right){\rm d}x}$
$=\left(\frac{\lambda_{0}+\eta_{0}\beta}{\lambda_{0}-\eta_{0}\beta}\right)^{\frac{d}{2}}\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\left\|\nabla V_{0}(x^{\prime})\right\|^{2}\right)$
$\leq\exp\left(\frac{\eta_{0}\beta d}{\lambda_{0}-\eta_{0}\beta}\right)\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\left\|\nabla V_{0}(x^{\prime})\right\|^{2}\right)$
@@
We choose $\lambda_{0}=\eta_{0}\beta d$ such that $\exp\left(\frac{\eta_{0}\beta d}{\lambda_{0}-\eta_{0}\beta}\right)\lesssim 1$. With this $\lambda_{0}$, $\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\left\|\nabla V_{0}(x^{\prime})\right\|^{2}\right)$ is also $O(1)$ as long as $\left\|\nabla V_{0}(x^{\prime})\right\|^{2}\lesssim\eta_{0}\beta d^{2}$.
@@
$\left\|\nabla V_{0}(x^{\prime})\right\|^{2}=\left\|\nabla V_{0}(x^{\prime})-\nabla V_{0}(x^{\prime\prime})\right\|^{2}\leq(\eta_{0}\beta d+\eta_{0}\beta)^{2}\left\|x^{\prime}-x^{\prime\prime}\right\|^{2}.$
Therefore, to guarantee $\left\|\nabla V_{0}(x^{\prime})\right\|^{2}\lesssim\eta_{0}\beta d^{2}$, it suffices to find an $x^{\prime}$ such that $\left\|x^{\prime}-x^{\prime\prime}\right\|\lesssim\frac{1}{\sqrt{\eta_{0}\beta}}$.
We first derive an upper bound of $\left\|x^{\prime\prime}\right\|$ under Assumption 1. Since $x_{*}$ is a global minimizer of $V$,
$\beta d\left\|x^{\prime\prime}\right\|$ $=\left\|\nabla V(x^{\prime\prime})\right\|=\left\|\nabla V(x^{\prime\prime})-\nabla V(x_{*})\right\|$
$\leq\beta\left\|x^{\prime\prime}-x_{*}\right\|\leq\beta(\left\|x^{\prime\prime}\right\|+R),$
$\implies\left\|x^{\prime\prime}\right\|$ $\leq\frac{R}{d-1}\lesssim\frac{R}{d}.$
When $R=0$, i.e., 00 is a known global minimizer of both $V$ and $V_{0}$, we can directly set $x^{\prime}=0$; otherwise, we need to run optimization algorithms to find such an $x^{\prime}$. According to standard results in convex optimization (see, e.g., [garrigos2023handbook, Theorem 3.6]), running gradient descent on function $V_{0}$ with step size $\frac{1}{\lambda_{0}+\eta_{0}\beta}$ yields the following convergence rate: starting from $x_{0}=0$, the $t$-th iterate $x_{t}$ satisfies
$\left\|x_{t}-x^{\prime\prime}\right\|^{2}\leq\left(1-\frac{\lambda_{0}-\eta_{0}\beta}{\lambda_{0}+\eta_{0}\beta}\right)^{t}\left\|0-x^{\prime\prime}\right\|^{2}\lesssim\left(\frac{2}{d}\right)^{t}\frac{R^{2}}{d^{2}}\leq\frac{R^{2}}{{\rm e}^{t}d^{2}},$
where the last inequality holds when $d\geq 6$. Thus, $\log\frac{\eta_{0}\beta R^{2}}{d^{2}}+O(1)$ iterations are sufficient to find a desired $x^{\prime}$. We summarize the rejection sampling algorithm in Algorithm 2. In conclusion, precisely sampling from $\pi_{0}$ requires
$O\left(1\vee\log\frac{\eta_{0}\beta R^{2}}{d^{2}}\right)=\widetilde{O}\left(1\right)$
calls to the oracle of $V$ and $\nabla V$ in expectation.
@@
1 Let $V_{0}:=\eta_{0}V+\frac{\lambda_{0}}{2}\left\|\cdot\right\|^{2}$;

### v2

@@ 7 Conclusions and Future Work
In this paper, we have explored the complexity of ALMC for sampling from a non-log-concave probability measure, circumventing the reliance on log-concavity or isoperimetric inequalities. Central to our analysis are the Girsanov theorem and optimal transport techniques, providing a novel approach. While our proof primarily focuses on the annealing scheme as described in Eq. 5, it can potentially be adapted to more general interpolations. Further exploration of these applications will be a key direction for future research. Technically, our proof methodology could be expanded to a broader range of target distributions beyond those with smooth potentials, such as those with Hölder-continuous gradients (Chatterji et al., 2020; Liang & Chen, 2022; Fan et al., 2023). Eliminating the assumption of global minimizers of the potential function would further enhance the practical applicability of our algorithm. Finally, while this work emphasizes the upper bounds for non-log-concave sampling, exploring the tightness of these bounds and investigating potential lower bounds for this problem (Chewi et al., 2023a; b) are intriguing avenues for future research.
Acknowledgments
We would like to thank Ye He, Sinho Chewi, Matthew S. Zhang, Omar Chehab, Adrien Vacher, and Anna Korba for insightful discussions. MT is supported by the National Science Foundation under Award No. DMS-1847802, and WG and YC are supported by the National Science Foundation under Award No. ECCS-1942523 and DMS-2206576.
Appendix A Proof of Results on Action
Lemma 4 (Sufficient Condition for Absolute Continuity (Informal)).
Assume that a curve of probability distributions $(\pi_{\theta})_{\theta\in[0,1]}$ on $\mathbb{R}^{d}$ has a density $\pi:[0,1]\times\mathbb{R}^{d}\ni(\theta,x)\mapsto\pi_{\theta}(x)>0$ that is jointly $C^{1}$. Then, this curve is AC.
Proof.
The following proof is informal. We leave the task of formalizing this statement for future work.
We first prove the result in one-dimension. Let $F_{\theta}(x)=\int_{-\infty}^{x}\pi_{\theta}(u){\rm d}u$ be the c.d.f. of $\pi_{\theta}$, which is strictly increasing as $\pi_{\theta}(x)>0$ everywhere. As a result, its inverse $F_{\theta}^{-1}$ is well-defined, and $F^{-1}_{\cdot}(\cdot)$ is also jointly $C^{1}$. We know from Ambrosio et al. (2008, Theorem 6.0.2) that
$W_{2}^{2}(\pi_{\theta},\pi_{\theta+\delta})=\int_{0}^{1}|F_{\theta+\delta}^{-1}(q)-F_{\theta}^{-1}(q)|^{2}{\rm d}q.$
The above quantity should be $O(\delta^{2})$ as $\delta\to 0$ due to the $C^{1}$ property, and hence the curve is AC.
In $d$-dimensions, consider the one-dimensional subspace robust $\text{W}_{\text{2}}$ distance (Paty & Cuturi, 2019) defined as
$S_{2}(\mu,\nu)=\sup_{v\in\mathbb{S}^{d-1}}W_{2}(v_{\sharp}\mu,v_{\sharp}\nu),$
where $\mathbb{S}^{d-1}=\{v\in\mathbb{R}^{d}:\|v\|=1\}$ is the set of unit vectors in $\mathbb{R}^{d}$, and $v_{\sharp}\mu$ (respectively, $v_{\sharp}\nu$) is the law of $\left\langle v,X\right\rangle$ when $X\sim\mu$ (respectively, $X\sim\nu$), which is a probability measure in $\mathbb{R}$. By Paty & Cuturi (2019, Proposition 2), $W_{2}(\mu,\nu)\leq\sqrt{d}S_{2}(\mu,\nu)$. Similar argument as above shows that $W_{2}^{2}(v_{\sharp}\pi_{\theta},v_{\sharp}\pi_{\theta+\delta})=O(\delta^{2})$, and hence $W_{2}^{2}(\pi_{\theta},\pi_{\theta+\delta})=O(\delta^{2})$. ∎
Proof of Lem. 3.
  1. 1.
    By Cauchy-Schwarz inequality,
    ${\cal A}=\int_{0}^{1}|\dot{\rho}|_{t}^{2}\mathrm{d}t\int_{0}^{1}1\mathrm{d}t\geq\left(\int_{0}^{1}|\dot{\rho}|_{t}\mathrm{d}t\right)^{2}.$
    It remains to prove that $W_{2}(\rho_{0},\rho_{1})\leq\int_{0}^{1}|\dot{\rho}|_{t}\mathrm{d}t$. In fact, take any sequence $10=t_{0}<t_{1}<...<t_{N}=$ 1, and let $\Delta t:=\max_{1\leq i\leq N}(t_{i}-t_{i-1})$. By triangle inequality of the $\text{W}_{\text{2}}$ distance,
    $W_{2}(\rho_{0},\rho_{1})$ $\leq\sum_{i=1}^{N}W_{2}(\rho_{t_{i}},\rho_{t_{i-1}})$
    $=\sum_{i=1}^{N}\frac{W_{2}(\rho_{t_{i}},\rho_{t_{i-1}})}{t_{i}-t_{i-1}}(t_{i}-t_{i-1})$
    $=\sum_{i=1}^{N}(|\dot{\rho}|_{t_{i}}+o(\Delta t))(t_{i}-t_{i-1})\to\int_{0}^{1}|\dot{\rho}|_{t}\mathrm{d}t.$
    Another derivation is by investigating the variational representations of these two quantities. First, by Lem. 2, one can write
    ${\cal A}=\int_{0}^{1}|\dot{\rho}|_{t}^{2}\mathrm{d}t=\inf_{v_{t}:~{}\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0}\int_{0}^{1}\|v_{t}\|_{L^{2}(\rho_{t})}^{2}{\rm d}t.$
    On the other hand, the Benamou-Brenier formula (Ambrosio et al., 2008, Equation 8.0.3) implies
    $W_{2}^{2}(\rho_{0},\rho_{1})=\inf_{(\rho_{t},v_{t}):~{}\partial_{t}\rho_{t}+\nabla\cdot(\rho_{t}v_{t})=0;~{}\rho_{t=0}=\rho_{0},\rho_{t=1}=\rho_{1}}\int_{0}^{1}\|v_{t}\|_{L^{2}(\rho_{t})}^{2}{\rm d}t.$
    Thus, one can easily conclude the desired inequality.
    To show that the inequality is attained when $(\rho_{t})_{t\in[0,1]}$ is the constant-speed Wasserstein geodesic, note that the derivation implies that the inequality is attained when $|\dot{\rho}|_{t}$ is a constant for all $t\in[0,1]$, and for any $0\leq t_{1}<t_{2}<t_{3}\leq 1$, $W_{2}(\rho_{t_{1}},\rho_{t_{2}})+W_{2}(\rho_{t_{2}},\rho_{t_{3}})=W_{2}(\rho_{t_{1}},\rho_{t_{3}})$. We refer readers to Ambrosio et al. (2008, Chapter 7.2) for the construction of the constant-speed Wasserstein geodesic.
  2. 2.
    It is known from Bakry et al. (2014) that $\pi$ satisfies $C$-LSI implies
    $\operatorname{KL}(\mu\|\pi)\leq\frac{C}{2}\mathbb{E}_{\mu}\left\|\nabla\log\frac{{\rm d}\mu}{{\rm d}\pi}\right\|^{2},~{}~{}~{}~{}\operatorname{KL}(\mu\|\pi)\geq\frac{1}{2C}W_{2}^{2}(\mu,\pi),~{}\forall\mu.$
    Therefore,
    $\frac{1}{\delta^{2}}W_{2}^{2}(\rho_{t+\delta},\rho_{t})\leq C_{\rm LSI}(\rho_{t})^{2}\int\frac{\|\nabla\log\rho_{t+\delta}-\nabla\log\rho_{t}\|^{2}}{\delta^{2}}{\rm d}\rho_{t+\delta}.$
    By letting $\delta\to 0$ and assuming regularity conditions, we have
    $|\dot{\rho}|_{t}^{2}\leq C_{\rm LSI}(\rho_{t})^{2}\int\|\partial_{t}\nabla\log\rho_{t}\|^{2}{\rm d}\rho_{t}=C_{\rm LSI}(\rho_{t})^{2}\|\partial_{t}\nabla\log\rho_{t}\|^{2}_{L^{2}(\rho_{t})}$
  3. 3.
    By Liu (2020), we know that $\pi$ satisfies $C$-PI implies
    $W_{2}^{2}(\mu,\pi)\leq 2C\left(\int\frac{{\rm d}\mu}{{\rm d}\pi}\mathrm{d}\mu-1\right),~{}\forall\mu.$
    Therefore,
    $\frac{1}{\delta^{2}}W_{2}^{2}(\rho_{t+\delta},\rho_{t})\leq 2C_{\rm PI}(\rho_{t})\int\frac{(\rho_{t+\delta}-\rho_{t})^{2}}{\delta^{2}\rho_{t}}{\rm d}x.$
    By letting $\delta\to 0$, we have
    $|\dot{\rho}|_{t}^{2}\leq 2C_{\rm PI}(\rho_{t})\int\frac{(\partial_{t}\rho_{t})^{2}}{\rho_{t}}{\rm d}x=2C_{\rm PI}(\rho_{t})\|\partial_{t}\log\rho_{t}\|_{L^{2}(\rho_{t})}^{2}.$
$\square$
Proof of Ex. 1.
  1. 1.
    The reparameterized curve $(p_{s})_{s\in[0,S]}$ satisfies the standard heat equation:
    $\partial_{s}p_{s}=\Delta p_{s}\implies\partial_{s}p_{s}+\nabla\cdot(p_{s}(-\nabla\log p_{s}))=0,$
    which means $(-\nabla\log p_{s})_{s\in[0,S]}$ generates $(p_{s})_{s\in[0,S]}$. According to the uniqueness result in Lem. 2 and also Ambrosio et al. (2008, Theorem 8.3.1), the Lebesgue-a.e. unique vector field $(v_{t}^{*})_{t\in[a,b]}$ that generates $(\rho_{t})_{t\in[a,b]}$ and satisfies $|\dot{\rho}|_{t}=\|v^{*}_{t}\|_{L^{2}(\rho_{t})}$ for Lebesgue-a.e. $t\in[a,b]$ can be written in a gradient field $v^{*}_{t}=\nabla\phi_{t}$ for some $\phi_{t}:\mathbb{R}^{d}\to\mathbb{R}$. This implies that
    $|\dot{p}|_{s}=\|\nabla\log p_{s}\|^{2}_{L^{2}(p_{s})}.$
    By the time change $[0,S]\ni s\mapsto\theta=s/S\in[0,1]$ and taking integral, we obtain the desired result.
  2. 2.
    Since $p_{s}=\sum_{i=1}^{N}w_{i}\operatorname{{\cal N}}\left(\mu_{i},(\sigma^{2}+2s)I\right)$, by Lem. 7, we know that $-\nabla^{2}\log p_{s}\preceq\frac{1}{\sigma^{2}+2s}I$. Thus Lem. 6 implies $\|\nabla\log p_{s}\|^{2}_{L^{2}(p_{s})}\leq\frac{d}{\sigma^{2}+2s}$, so
    ${\cal A}\leq S\int_{0}^{S}\frac{d}{\sigma^{2}+2s}\mathrm{d}s=\frac{Sd}{2}\log\left(1+\frac{2S}{\sigma^{2}}\right).$
    Finally, the LSI and PI constant of mixture of Gaussian distributions are, in general, exponential with respect to the maximum distance between the means. To illustrate this, consider a simpler case where there are only 2 Gaussian components: $\rho=\frac{1}{2}{\cal N}(0,\sigma^{2}I)+\frac{1}{2}{\cal N}(y,\sigma^{2}I)$. Schlichting (2019, Section 4.1) has shown that $C_{\rm PI}(\rho)\leq C_{\rm LSI}(\rho)\leq\frac{\sigma^{2}}{2}\left({\rm e}^{\|y\|^{2}/\sigma^{2}}+3\right)$, while Dong & Tong (2022, Proposition 1) has shown that when $\|y\|\gtrsim\sigma\sqrt{d}$, $C_{\rm LSI}(\rho)\geq C_{\rm PI}(\rho)\gtrsim\frac{\sigma^{4}}{\|y\|^{2}}{\rm e}^{\Omega(\|y\|^{2}/\sigma^{2})}$. $\square$
Appendix B Sampling from $\pi_{0}$
Lemma 5.
When $\eta_{0}=0$, $\pi_{0}=\operatorname{{\cal N}}\left(0,\lambda_{0}^{-1}I\right)$ can be sampled directly; when $\eta_{0}\in(0,1]$, we choose $\lambda_{0}=\eta_{0}d\beta$, so that under Assump. 2, it takes $O\left(1\vee\log\frac{\eta_{0}\beta R^{2}}{d^{2}}\right)=\widetilde{O}(1)$ queries to the oracle of $V$ and $\nabla V$ in expectation to precisely sample from $\pi_{0}$ via rejection sampling.
Proof.
@@ Appendix B Sampling from $\pi_{0}$
Our proof is inspired by Liang & Chen (2023). We only consider the nontrivial case $\eta_{0}\in(0,1]$. The potential of $\pi_{0}$ is $V_{0}:=\eta_{0}V+\frac{\lambda_{0}}{2}\|\cdot\|^{2}$, which is $(\lambda_{0}-\eta_{0}\beta)$-strongly-convex and $(\lambda_{0}+\eta_{0}\beta)$-smooth. Note that for any fixed point $x^{\prime}\in\mathbb{R}^{d}$,
@@
$\pi_{0}(x)\propto\exp\left(-V_{0}(x)\right)\leq\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right).$
@@
The r.h.s. is the unnormalized density of $\pi_{0}^{\prime}:=\operatorname{{\cal N}}\left(x^{\prime}-\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}-\eta_{0}\beta},\frac{1}{\lambda_{0}-\eta_{0}\beta}I\right)$. The rejection sampling algorithm is as follows: sample $X\sim\pi_{0}^{\prime}$, and accept $X$ as a sample from $\pi_{0}$ with probability
@@
$\exp\left(-V_{0}(X)+V_{0}(x^{\prime})+\left\langle\nabla V_{0}(x^{\prime}),X-x^{\prime}\right\rangle+\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right)\in(0,1].$
@@
By Chewi (2024, Theorem 7.1.1), the number of queries to the oracle of $V$ until acceptance follows a geometric distribution with mean
@@
$\frac{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right)\mathrm{d}x}{\int\exp\left(-V_{0}(x)\right)\mathrm{d}x}$
@@
$\leq\frac{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right)\mathrm{d}x}{\int\exp\left(-V_{0}(x^{\prime})-\left\langle\nabla V_{0}(x^{\prime}),x-x^{\prime}\right\rangle-\frac{\lambda_{0}+\eta_{0}\beta}{2}\|x-x^{\prime}\|^{2}\right)\mathrm{d}x}$
@@
$=\frac{\int\exp\left(-\left\langle\nabla V_{0}(x^{\prime}),x\right\rangle-\frac{\lambda_{0}-\eta_{0}\beta}{2}\|x\|^{2}\right)\mathrm{d}x}{\int\exp\left(-\left\langle\nabla V_{0}(x^{\prime}),x\right\rangle-\frac{\lambda_{0}+\eta_{0}\beta}{2}\|x\|^{2}\right)\mathrm{d}x}$
@@
$=\frac{\int\exp\left(-\frac{\lambda_{0}-\eta_{0}\beta}{2}\left\|x+\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}-\eta_{0}\beta}\right\|^{2}+\frac{\|\nabla V_{0}(x^{\prime})\|^{2}}{2(\lambda_{0}-\eta_{0}\beta)}\right)\mathrm{d}x}{\int\exp\left(-\frac{\lambda_{0}+\eta_{0}\beta}{2}\left\|x+\frac{\nabla V_{0}(x^{\prime})}{\lambda_{0}+\eta_{0}\beta}\right\|^{2}+\frac{\|\nabla V_{0}(x^{\prime})\|^{2}}{2(\lambda_{0}+\eta_{0}\beta)}\right)\mathrm{d}x}$
$=\left(\frac{\lambda_{0}+\eta_{0}\beta}{\lambda_{0}-\eta_{0}\beta}\right)^{\frac{d}{2}}\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\|\nabla V_{0}(x^{\prime})\|^{2}\right)$
$\leq\exp\left(\frac{\eta_{0}\beta d}{\lambda_{0}-\eta_{0}\beta}\right)\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\|\nabla V_{0}(x^{\prime})\|^{2}\right)$
@@
We choose $\lambda_{0}=\eta_{0}\beta d$ such that $\exp\left(\frac{\eta_{0}\beta d}{\lambda_{0}-\eta_{0}\beta}\right)\lesssim 1$. With this $\lambda_{0}$, $\exp\left(\frac{\eta_{0}\beta}{\lambda_{0}^{2}-\eta_{0}^{2}\beta^{2}}\|\nabla V_{0}(x^{\prime})\|^{2}\right)$ is also $O(1)$ as long as $\|\nabla V_{0}(x^{\prime})\|^{2}\lesssim\eta_{0}\beta d^{2}$.
@@
$\|\nabla V_{0}(x^{\prime})\|^{2}=\|\nabla V_{0}(x^{\prime})-\nabla V_{0}(x^{\prime\prime})\|^{2}\leq(\eta_{0}\beta d+\eta_{0}\beta)^{2}\|x^{\prime}-x^{\prime\prime}\|^{2}.$
Therefore, to guarantee $\|\nabla V_{0}(x^{\prime})\|^{2}\lesssim\eta_{0}\beta d^{2}$, it suffices to find an $x^{\prime}$ such that $\|x^{\prime}-x^{\prime\prime}\|\lesssim\frac{1}{\sqrt{\eta_{0}\beta}}$.
We first derive an upper bound of $\|x^{\prime\prime}\|$ under Assump. 2. Since $x_{*}$ is a global minimizer of $V$,
$\beta d\|x^{\prime\prime}\|$ $=\|\nabla V(x^{\prime\prime})\|=\|\nabla V(x^{\prime\prime})-\nabla V(x_{*})\|$
$\leq\beta\|x^{\prime\prime}-x_{*}\|\leq\beta(\|x^{\prime\prime}\|+R),$
$\implies\|x^{\prime\prime}\|$ $\leq\frac{R}{d-1}\lesssim\frac{R}{d}.$
When $R=0$, i.e., 00 is a known global minimizer of both $V$ and $V_{0}$, we can directly set $x^{\prime}=0$; otherwise, we need to run optimization algorithms to find such an $x^{\prime}$. According to standard results in convex optimization (see, e.g., Garrigos & Gower (2023, Theorem 3.6)), running gradient descent on function $V_{0}$ with step size $\frac{1}{\lambda_{0}+\eta_{0}\beta}$ yields the following convergence rate: starting from $x_{0}=0$, the $t$-th iterate $x_{t}$ satisfies
$\|x_{t}-x^{\prime\prime}\|^{2}\leq\left(1-\frac{\lambda_{0}-\eta_{0}\beta}{\lambda_{0}+\eta_{0}\beta}\right)^{t}\|0-x^{\prime\prime}\|^{2}\lesssim\left(\frac{2}{d}\right)^{t}\frac{R^{2}}{d^{2}}\leq\frac{R^{2}}{{\rm e}^{t}d^{2}},$
where the last inequality holds when $d\geq 6$. Thus, $\log\frac{\eta_{0}\beta R^{2}}{d^{2}}+O(1)$ iterations are sufficient to find a desired $x^{\prime}$. We summarize the rejection sampling in Alg. 2. In conclusion, precisely sampling from $\pi_{0}$ requires
$O\left(1\vee\log\frac{\eta_{0}\beta R^{2}}{d^{2}}\right)=\widetilde{O}(1)$
calls to the oracle of $V$ and $\nabla V$ in expectation. ∎
@@
1 Let $V_{0}:=\eta_{0}V+\frac{\lambda_{0}}{2}\|\cdot\|^{2}$;